OpenStudy (anonymous):

Mary is 14 years younger than her sister Jane. In four years, Jane will be 3 times older than Mary. How old is Jane now? 3 7 17 21 Is she 21

OpenStudy (tyteen4a03):

Shortcut for verification: Let the age of Jane be 21, then follow the instructions in the question. If she's the 21, the outcome of whatever is in there will be consistent with Mary's age.

OpenStudy (anonymous):

This requires 2 equations: one to describe their current ages and one to describe their ages in 4 years.

OpenStudy (anonymous):

no its 17 what m getting using 2 equations where is assume marry as m and jane as j 1st eqn m= j-14 2nd eqn j+4=3*(m+4) solve and get answers

OpenStudy (e.cociuba):

Present age of Mary=x, Jane=y x=y-14 and y+4=3(x+4) Solving we get,x=3,y=17 After 4 more years Jane will be 21, so wait for 4 more years.

OpenStudy (anonymous):

Those equations are correct. Now use substitution to solve them.

OpenStudy (anonymous):

Since you are looking for Jane's age, I would use \(m=j-14\) which is already solved for m and substitute it into \(j+4=3(m+4)\).

OpenStudy (anonymous):

By the way, is the "sigma Equation" button not working for anyone else?

OpenStudy (anonymous):

It seems to be working... \[\sum_{a}^{b}\]

OpenStudy (anonymous):

Thank you. At least the window comes up for me.